OpenStudy (anonymous):

write an expression for this $90 Includes unlimited minutes

OpenStudy (anonymous):

We need more information to help you out.

OpenStudy (anonymous):

$40/mo, incl. 300 min + $0.40 per extra minute. Bill was $168 168 = 40 + 0.40a (again, a = number of minutes paid extra) 168 - 40 = 0.40a 128 = 0.40a a = 128 / 0.40 = 320 additional minutes

OpenStudy (anonymous):

Apparently not, good job Dakota.

OpenStudy (anonymous):

anymore questions @iceicebaby

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

$15 monthly fee and $0.45 per minute and expression for this to